You have an internship in a local pharmacy. A rece
Answer #1

This is dilution of the 10% mix.. with 5% mix...

so...

total volume --> 250 mL

Final concentration --> 8.5 %

so..

0.05*v1 + 0.10*v2 = 0.085*v3

note that v3 = 250 so

0.05*v1 + 0.10*v2 = 0.085*250

and we know that

v1 + v2 = 250

so.. v2 = 250-v1

0.05*v1 + 0.10*(250-v1) = 0.085*250

0.05*v1 + 25 - 0.10 v1 = 21.25

v1(0.05-0.1) = 21.25-25

V1 = (21.25-25)/(0.05-0.1)

V1 = 75 mL of mix 1 required (5%)

V2 = 250-75= 175 mL of 10% mix required